A fire cracker tossed in the air at 60 degree with horizontal at a speed of30.0 m/s exploded 15m away. (A) how long was the fire cracker in the air before it exploded? (B) how high did it go?
1
Expert's answer
2020-11-10T06:56:24-0500

a)


"t=\frac{2v\sin{60}}{g}=\frac{2(30)\sin{60}}{9.8}=5.3\ s"

b)


"H=\frac{v^2\sin^2{60}}{2g}=\frac{(30)^2\sin^2{60}}{2\cdot9.8}=34.4\ s"